OpenStudy (anonymous):

Goal: to demonstrate the use of equations and inequalities in a real-world setting Role: You are the marketing manager at a cell phone provider. Audience: Customers who are thinking of changing their cell phone plan Situation: You have been asked to write a report for customer who is having a difficult time determining the best cell phone plan for their needs. Performance Task: Your company offers three cell phone plans. The details are below: Plan A allows 450 minutes for $39.99 with .25/minute for each minute over 450 Plan B allows 900 minutes for $59.99 with .30/minut

OpenStudy (anonymous):

Someone/Anyone PLEASE HELP

OpenStudy (anonymous):

Can anyone solve ??

OpenStudy (ganpat):

wait, m trying..

OpenStudy (ganpat):

Equations, Plan A , 39.99 $ ; 450 + 0.25 (x) Plan B, 59.99$ ; 900 + 0.30 (x) Where x is number of minutes, exceeding the base rate.. so say, you talk for 1000 minutes, So as per plan A; 39.99 + 0.25 (1000-450) =39.99 + 137.5 = 177.49$ and as per plan B, 59.99 + 0.30 (1000-900) = 59.99 + 30 = 89.99 $ , 450 + 0.25 (x) < 900 + 0.30 (x)...

OpenStudy (ganpat):

this is best, what i can get.. hope u understand that..